OpenStudy (anonymous):

What is the length of the minor axis of the ellipse given by the equation x^2/144 + y^2/25 = 1?

OpenStudy (mertsj):

\[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\]

OpenStudy (mertsj):

That is the equation of an ellipse like yours. The length of the minor axis is 2b

OpenStudy (anonymous):

So it would 5x2. Which would be 10.

OpenStudy (mertsj):

yep